Write down energy flow equation for non flow process. Estimate heat flow to/from the system.


Specific internal energy is 420kJ/kg at the beginning of the expansion and specific internal energy is


200kJ/kg at the end of the process. The work done by the system is 100kJ/kg during the process.

1
Expert's answer
2021-11-08T17:12:01-0500

"U_1=420KJ/kg\\U_2=200KJ/kg"


"\u2206U=U_2-U_1=420-200=220KJ\/kg"

"\frac{∆U}{m}=s∆T"

"s\u2206T=220KJ\/kg"